OpenStudy (cosma09):

Over a two-hour time period, a snail moved 108 inches. How far is this in yards?

OpenStudy (hockeychick23):

12 inches in a foot, 3 feet in a yard 108/12= 9 feet 9/3= 3 yards

OpenStudy (cosma09):

It takes 24 cups of chicken broth to make a chicken soup recipe. How much is this in quarts?

OpenStudy (hockeychick23):

1 quart= 4 cups 24/4= 6 quarts

OpenStudy (cosma09):

Click on "Explain" if you need help. Three pumps are pumping fluid at constant rates. For each, the amount of fluid pumped depends on time, as shown below. Pump A Pump B Pump C x1234y601201802403003604204800 Amountoffluid(L) Time(hr) Time (hours) 0 1 2 3 Amount of fluid (liters) 0 90 180 270 =f110t f is the amount of fluid in liters t is the time in hours Answer the questions below about the pumps. (a) What is the rate of each pump? PumpA:liters/hour PumpB:liters/hour PumpC:liters/hour (b) Which pump is the slowest? Pump A Pump B Pump C (c) How much fluid does each one pump in 3 hours? PumpA:liters PumpB:liters PumpC:liters ClearUndoHelp Next >> Explain